Students should always be careful of the chemical reaction between jewelry and laboratory reagents. The table below shows the reduction reactions and potentials for some common lab reagents and metals.

Ox + ne1- Red Eo(V)
Ni2+(aq) + 2e1- ↔ Ni(s) -0.23
SO42-(aq) + 4H1+(aq) + 2e1- ↔H2SO3(aq) + H2O +0.20
Cu2+(aq) + 2e1- ↔ Cu(s) +0.34
Ag1+(aq) + e1- ↔ Ag(s) +0.80
Pt2+(aq) + 2e1- ↔ Pt(s) +1.19
Cr2O72-(aq) + 14H1+(aq) + 6e1- ↔ 2Cr3+(aq) + 7H2O +1.33
Au3+(aq) + 3e1- ↔ Au(s) +1.50
MnO41-(aq) + 8H1+(aq) + 5e1- ↔ Mn2+(aq) + 4H2O +1.51

(a) What is the net redox reaction that occurs when Au comes into contact with acidic KMnO4? (Use the lowest possible coefficients. Omit states-of-matter from your answer.)

(b) What is E°cell for the reaction?

________ V

(c) Which metal is the least reactive?

Ag

Au   

Cu

Ni

Pt


(d) Which reagent is the most reactive?

Cr2O72−

MnO4   

SO42−
Answer #1

a)

oxidation reaction (Anode)

Au - 3e  \to Au3+ (1) Ered0 = + 1.50 V

reduction reaction (Cathode)

MnO4- + 8H+ + 5e  \to Mn2+ + 4 H2O (2) E0red = + 1.51 V  

Eq. 1 *5 + Eq.2*3

Balanced reaction is

5 Au + 3 MnO4- + 24 H+\to 3 Mn2+ + 5 Au3+ + 12 H2O

b)

E0cell = E0cathode - E0anode = 1.51 - 1.50 = +  0.01 V

C)

Least reactive metal has highest value of standard reduction potential.

Au has highest value standard reduction potential among the given metals . So Au is least reactive.

d)

Standard reduction potential of MnO4- is more positive than Cr2O72- and SO42- hence MnO4- is stronger oxidizing agent hence most reactive.